# Synesthesia of Machines (SoM)-Enhanced Wideband Multi-User CSI Learning With LiDAR Sensing

### v2

@@ I Introduction
In conclusion, LE-CLN stands out as an innovative channel estimator tailored for wideband multi-user MIMO-OFDM systems, effectively utilizing the inherent correlation between LiDAR environmental detection results and CSI. Through simulations, LE-CLN demonstrates substantial enhancements in channel estimation accuracy by integrating out-of-band wireless channel insights obtained from LiDAR, particularly in scenarios with limited pilot signals.
@@ II System Model
We consider a time division duplex (TDD) based multi-user wideband hybrid MIMO-OFDM communication system where a BS equipped with a uniform linear array uses $N_{\rm{s}}$ subcarriers to serve $K$ single-antenna users, as shown in Fig. 1. The BS is equipped with $N_{\rm{RF}}$ RF chains and $N_{\rm{t}}$ antennas (generally $N_{\rm{RF}}<N_{\rm{t}}$). The digital precoding matrix and analog precoding matrix adopted by BS are denoted by $\mathbf{F}_{\rm{B}}\in\mathbb{C}^{N_{\rm{RF}}\times N_{\rm{RF}}}$ and $\mathbf{F}_{\rm{R}}\in\mathbb{C}^{N_{\rm{t}}\times N_{\rm{RF}}}$ with $\|\mathbf{F}_{\rm{B}}\mathbf{F}_{\rm{R}}\|_{\rm{F}}^{2}=N_{\rm{RF}}$. $\mathbf{F}_{\rm{B}}$ is set as an identity matrix $\mathbf{I}_{N_{\rm{RF}}}$ similar to [8, 15] and is omitted in the following contents. Leveraging the channel reciprocity, the downlink channel can be directly derived through estimating the uplink channel.
In the considered system, multiple users adopt orthogonal subcarriers to transmit pilots to BS. The BS then estimates the channels of these subcarriers, allowing for the recovery of full downlink channels. Let $\mathbb{V}_{k}=\{v_{1},v_{2},\cdots,v_{K_{\rm{P}}}\}$ represent the set of subcarrier indices used by the $k$-th user for pilot transmission, where the subcarrier interval $\Delta_{v}=v_{i+1}-v_{i}=\lfloor\frac{N_{\rm{s}}}{K_{\rm{P}}}\rfloor$. It follows $\mathbb{V}_{i}\cap\mathbb{V}_{j}=\varnothing$ for $i\neq j$.
For the $k$-th user, the pilot symbols transmitted are denoted by $\mathbf{S}^{\rm{P}}_{k}\in\mathbb{C}^{K_{\rm{P}}\times N_{\rm{P}}}$ satisfying $\mathbb{E}\{\mathbf{S}^{\rm{P}}_{k}(\mathbf{S}^{\rm{P}}_{k})^{H}\}=\mathbf{I}_{K_{\rm{P}}}$, where $K_{\rm{P}}$ and $N_{\rm{P}}$ represent the number of pilots placed along the subcarrier and time-slot axes. The full downlink channel between the BS and the $k$-th user is denoted by $\mathbf{H}_{k}\triangleq[\mathbf{h}_{k}^{1},\mathbf{h}_{k}^{2},\cdots,\mathbf{h}_{k}^{N_{\rm{s}}}]\in\mathbb{C}^{N_{\rm{t}}\times N_{\rm{s}}}$, where $\mathbf{h}_{k}^{m}\in\mathbb{C}^{N_{\rm{t}}\times 1}$ is the channel over the $m$-th subcarrier for $m=1,2,\cdots,N_{\rm{s}}$. After the pilot transmission, the overall measurements $\mathbf{Y}^{\rm{P}}_{k}\in\mathbb{C}^{N_{\rm{RF}}\times N_{\rm{P}}}$ obtained by BS can be expressed as
$\mathbf{Y}^{\rm{P}}_{k}=\mathbf{F}_{\rm{R}}^{H}\mathbf{H}^{\rm{P}}_{k}\mathbf{S}^{\rm{P}}_{k}+\mathbf{F}_{\rm{R}}^{H}\mathbf{N}_{k},$ (1)
@@ III-A Coarse LiDAR Processing
Step 1: Raw Data Filtering. We first identify the ground points and eliminate them from the entire point cloud though they may contain multipath information from ground reflections. This is because the NN can hardly acquire such information due to its complex generation process. Therefore, this data filtering approach will alleviate the computational burden of NNs without excluding useful feature information.
@@ III-B CSI Feature Extraction via Coarse LiDAR Processing
where $N_{\rm{w}}$ determines the size of over-sampled subspace; $\epsilon=\frac{2\pi}{2N_{\rm{t}}}$ is the sampling interval of a uniform over-complete DFT codebook. The overall angle grid distribution $\mathbf{\Phi}$ is given by:
@@
where $\phi_{\rm{g}_{1}}$ and $\phi_{\rm{g}_{2}}$ are the end points of the plain angle subspace. Note that the angle grid within the over-sampled and plain subspace is uniformly partitioned, with an angle resolution of $\epsilon_{\rm{o}}=\frac{\epsilon}{2}$ in the over-complete subspace and $\epsilon_{\rm{g}}=\frac{2\pi-N_{\rm{o}}\epsilon_{\rm{o}}}{D-N_{\rm{o}}}$ in the plain subspace. $N_{\rm{o}}$ is the number of angle grids in the over-sampled subspace, which is determined by $N_{\rm{w}}$. For instance, considering the second distribution case in Eq. (3) as a reference, the angle grid $\phi_{\emph{i}}$ in different subspaces can be calculated as:
$\phi_{\emph{i}}=\left\{\begin{array}[]{cl}(i-1)\epsilon_{\rm{g}},~{}\emph{i}\in[1,{\rm{o}_{1}})&\\ \phi_{\rm{o}_{1}}+(i-\rm{o}_{1})\epsilon_{\rm{o}},~{}\emph{i}\in[{\rm{o}_{1}},{\rm{o}_{2}}]&\\ \phi_{\rm{o}_{2}}+(i-\rm{o}_{2})\epsilon_{\rm{g}},~{}\emph{i}\in(\rm{o}_{2},\emph{D})&.\end{array}\right.$ (4)
@@
Based on $\mathbf{A}$, the antenna-domain channel can be transformed into angular domain by $\mathbf{H}^{\rm{a}}_{k}=\mathbf{A}^{H}\mathbf{H}_{k}$, where $\mathbf{H}^{\rm{a}}_{k}\in\mathbb{C}^{D\times K_{\rm{P}}}$ is the angular-domain channel between the $k$-th user and BS of the pilot position. Using $(\mathbf{A}^{H})^{\dagger}\mathbf{A}^{H}=\mathbf{I}$, the received pilot signals can be rewritten as:
$\mathbf{Y}^{\rm{P}}_{k}=\underbrace{\mathbf{F}_{\rm{R}}^{H}(\mathbf{A}^{H})^{\dagger}}_{\Theta}\mathbf{H}^{\rm{a}}_{k}\mathbf{S}^{\rm{P}}_{k}+\mathbf{F}_{\rm{R}}^{H}\mathbf{N}^{\rm{P}}_{k}.$ (5)
Here, $\Theta=\mathbf{F}_{\rm{R}}^{H}(\mathbf{A}^{H})^{\dagger}$ is defined as the sensing matrix. The pilot observations $\mathbf{Y}^{\rm{P}}_{k}$ are projected into the sensing matrix space by $\tilde{\mathbf{Y}}^{\rm{P}}_{k}=\Theta^{\dagger}\mathbf{Y}^{\rm{P}}_{k}$, which captures user-localized angular features. The complex value $\tilde{\mathbf{Y}}^{\rm{P}}_{k}\in\mathbb{C}^{D\times N_{\rm{p}}}$ is then transformed into real value input $\tilde{\mathbf{Y}}^{\rm{r}}_{k}\in\mathbb{R}^{2\times D\times N_{\rm{p}}}$ by treating the imaginary and real parts as two parallel dimensions. $\tilde{\mathbf{Y}}^{\rm{r}}_{k}$ serves as the input for a customized two-layer CNN, referred to as the pilot-based CSI feature extraction CNN (PCF-CNN).
@@ III-C OOB CSI Feature Extraction via Fine LiDAR Processing
In Eq. (6), $f=f_{\rm{up}}+f_{\rm{down}}$ is the vertical field of view (FoV) of LiDAR; $f_{\rm{up}}$ and $f_{\rm{down}}$ are the upward and downward FoV, respectively. The width ($w$) and height ($h$) of range image are determined by the horizontal angular resolution $a_{\rm{h}}$ and the number of beams $M_{\rm{b}}$ of the LiDAR, such that $w=\frac{360^{\circ}}{a_{\rm{h}}}$ and $h=M_{\rm{b}}$.
Step 2: Multiple Signal Propagation (SP)-Feature Channels Superposition. Based on the LiDAR range image $\mathbf{R}_{\rm{D}}$, we propose to construct a multi-SP-feature superposed range image $\mathbf{R}_{\rm{D}}^{\rm{RF}}\in\mathbb{R}^{3\times w\times h}$. The first channel represents the original range image, denoted as $\mathbf{R}_{\rm{D}}^{\rm{RF}}[1,u_{i},v_{i}]=r_{i}$ where $r_{i}$ is the range at the pixel $\mathbf{r}_{i}$. The second channel is referred to as the equivalent small-scale fading map, which aims to provide information about potential multipath components. The pixel values in this channel, $\mathbf{R}_{\rm{D}}^{\rm{RF}}[2,u_{i},v_{i}]$, can take one of three possible values: receiver, vehicles near the receiver (potential scatterers), and other objects. The third channel is known as the equivalent large-scale fading map, designed to offer insights into signal amplitude variation characteristics. The pixel value in this channel, $\mathbf{R}_{\rm{D}}^{\rm{RF}}[3,u_{i},v_{i}]$, represents the path loss of the signal transmitted from transmitter to the point $\mathbf{r}_{i}$. The pixel values of the second and third channels are determined through specific criteria and considerations related to the environmental detection results acquired from LiDAR, and are expressed as:
$\mathbf{R}_{\rm{D}}^{\rm{RF}}[2,u_{i},v_{i}]=\left\{\begin{array}[]{cl}2&,~{}\mathbf{p}_{i}\in\mathcal{V}_{\rm{r}}\\ 1&,~{}\mathbf{p}_{i}\in\mathcal{V}_{\rm{s}}\\ 0&,~{}\text{otherwise}\end{array}\right.$ (7a)
@@ III-D Feature Fusion via Attention
Although OOB wireless features closely related to CSI have been extracted from LiDAR and fine-grained user-localized CSI features have been extracted from pilots, a key challenge that remains to be addressed is how to intelligently regulate the utilization of both modalities under varying channel conditions. Drawing inspiration from the SENet channel attention mechanism [17], we have designed an adaptive feature weight control (AFWC) module to facilitate element-wise weight allocation between CSI features from the two modalities. Denoting the CSI features extracted from pilots and LiDAR as $\mathbf{q}_{\rm{P}}$ and $\mathbf{q}_{\rm{L}}$, respectively, they are concatenated and input into the AFWC module, which learns the importance of different features and outputs the weight vector $\mathbf{w}$:
@@
Here, $\varphi_{\rm w}^{l}(\cdot)=f_{\rm{ReLU}}(\mathbf{W}_{\rm w}^{l}\mathbf{X}+\mathbf{b}_{\rm w}^{l}),l=1,2$ is the non-linear functions of the first and second fully connected (FC) layers with ReLU as the activation function; $\varphi_{\rm{w}}^{3}(\cdot)=f_{\rm{Sigmoid}}(\mathbf{W}_{\rm w}^{3}\mathbf{X}+\mathbf{b}_{\rm w}^{3})$ is the non-linear function of the third layer with Sigmoid as the activation function; $\oplus$ represents the concatenation operation. $\mathbf{X}$ denotes a certain input, while $\mathbf{W}_{\rm w}^{l}$ and $\mathbf{b}_{\rm w}^{l}$ denote the weight and bias of the $l$-th layer. Next, a weighted feature $\mathbf{q}^{\rm{w}}$ is formed by performing an element-wise product between the weight vector $\mathbf{w}$ and the raw multi-modal CSI feature $\mathbf{q}$, denoted as $\mathbf{q}^{\rm{w}}=\mathbf{q}\otimes\mathbf{w}$.
@@ III-E CSI Reconstruction
$\hat{\mathbf{H}}_{k}^{\rm{P}}=\text{MLP}_{\rm P}(\mathbf{q}^{\rm{w}},\mathbf{\Theta}_{\rm P})=\varphi_{\rm P}^{L_{\rm P}}(\varphi_{\rm P}^{L_{\rm P}-1}(\cdots\varphi_{\rm P}^{1}(\mathbf{q}^{\rm{w}})\cdots)).$ (9)

### v3

@@ I Introduction
In conclusion, LE-CLN stands out as an innovative channel estimator tailored for wideband multi-user MIMO-OFDM systems, effectively utilizing the inherent correlation between LiDAR environmental detection results and CSI. Through simulations, LE-CLN demonstrates substantial enhancements in channel estimation accuracy by integrating out-of-band wireless channel insights obtained from LiDAR, particularly in scenarios with limited pilot signals. [^1] Simulation codes are provided to reproduce the results presented in this paper: https://github.com/Haotian-Zhang-PKU/CSI-Learning-LECLN
@@ II System Model
We consider a time division duplex (TDD) based multi-user wideband hybrid MIMO-OFDM communication system where a BS equipped with a uniform linear array uses $N_{\rm{s}}$ subcarriers to serve $K$ single-antenna users, as shown in Fig. 1. The BS is equipped with $N_{\rm{RF}}$ RF chains and $N_{\rm{t}}$ antennas (generally $N_{\rm{RF}}<N_{\rm{t}}$). The digital precoding matrix and analog precoding matrix adopted by BS are denoted by $\mathbf{F}_{\rm{B}}\in\mathbb{C}^{N_{\rm{RF}}\times N_{\rm{RF}}}$ and $\mathbf{F}_{\rm{R}}\in\mathbb{C}^{N_{\rm{t}}\times N_{\rm{RF}}}$ with $\|\mathbf{F}_{\rm{R}}\mathbf{F}_{\rm{B}}\|_{\rm{F}}^{2}=N_{\rm{RF}}$. $\mathbf{F}_{\rm{B}}$ is set as an identity matrix $\mathbf{I}_{N_{\rm{RF}}}$ similar to [8, 15] and is omitted in the following contents. Leveraging the channel reciprocity, the downlink channel can be directly derived through estimating the uplink channel.
In the considered system, multiple users adopt orthogonal subcarriers to transmit pilots to BS. The BS then estimates the channels of these subcarriers, allowing for the recovery of full downlink channels. Let $\mathbb{V}_{k}=\{v_{1},v_{2},\cdots,v_{K_{\rm{P}}}\}$ represent the set of subcarrier indices used by the $k$-th user for pilot transmission, where the subcarrier interval $\Delta_{v}=v_{i+1}-v_{i}=\lfloor{N_{\rm{s}}}/{K_{\rm{P}}}\rfloor$. It follows $\mathbb{V}_{i}\cap\mathbb{V}_{j}=\varnothing$ for $i\neq j$.
For the $k$-th user, the pilot symbols transmitted are denoted by $\mathbf{S}^{\rm{P}}_{k}\in\mathbb{C}^{K_{\rm{P}}\times N_{\rm{P}}}$ satisfying $\mathbb{E}\{\mathbf{S}^{\rm{P}}_{k}(\mathbf{S}^{\rm{P}}_{k})^{\rm H}\}=\mathbf{I}_{K_{\rm{P}}}$, where $K_{\rm{P}}$ and $N_{\rm{P}}$ represent the number of pilots placed along the subcarrier and time-slot axes. The full downlink channel between the BS and the $k$-th user is denoted by $\mathbf{H}_{k}\triangleq[\mathbf{h}_{k}^{1},\mathbf{h}_{k}^{2},\cdots,\mathbf{h}_{k}^{N_{\rm{s}}}]\in\mathbb{C}^{N_{\rm{t}}\times N_{\rm{s}}}$, where $\mathbf{h}_{k}^{m}\in\mathbb{C}^{N_{\rm{t}}\times 1}$ is the channel over the $m$-th subcarrier for $m=1,2,\cdots,N_{\rm{s}}$. After the pilot transmission, the overall measurements $\mathbf{Y}^{\rm{P}}_{k}\in\mathbb{C}^{N_{\rm{RF}}\times N_{\rm{P}}}$ obtained by BS can be expressed as
$\mathbf{Y}^{\rm{P}}_{k}=\mathbf{F}_{\rm{R}}^{\rm H}\mathbf{H}^{\rm{P}}_{k}\mathbf{S}^{\rm{P}}_{k}+\mathbf{F}_{\rm{R}}^{\rm H}\mathbf{N}_{k},$ (1)
@@ III-A Coarse LiDAR Processing
Step 1: Raw Data Filtering. We first identify the ground points and eliminate them from the entire point cloud though they may contain multipath information from ground reflections. This is because the NN can hardly acquire such information due to its complex generation process. The data filtering procedure will alleviate the computational burden while retaining the primary useful features. It is worth noting that in cases where multipath components from ground reflections are non-negligible, more pilots will be needed, to supply features that cannot be acquired through LiDAR.
@@ III-B CSI Feature Extraction via Coarse LiDAR Processing
where $N_{\rm{w}}$ determines the size of over-sampled subspace; $\epsilon={2\pi}/{2N_{\rm{t}}}$ is the sampling interval of a uniform over-complete DFT codebook. The overall angle grid distribution $\mathbf{\Phi}$ is given by:
@@
where $\phi_{\rm{g}_{1}}$ and $\phi_{\rm{g}_{2}}$ are the end points of the plain angle subspace. Note that the angle grid within the over-sampled and plain subspace is uniformly partitioned, with an angle resolution of $\epsilon_{\rm{o}}={\epsilon}/{2}$ in the over-complete subspace and $\epsilon_{\rm{g}}=({2\pi-N_{\rm{o}}\epsilon_{\rm{o}}})/({D-N_{\rm{o}}})$ in the plain subspace. $N_{\rm{o}}$ is the number of angle grids in the over-sampled subspace, which is determined by $N_{\rm{w}}$. For instance, considering the second distribution case in Eq. (3) as a reference, the angle grid $\phi_{\emph{i}}$ in different subspaces can be calculated as:
$\phi_{\emph{i}}=\left\{\begin{array}[]{cl}(i-1)\epsilon_{\rm{g}},~{}\emph{i}\in[1,{\rm{o}_{1}})&\\ \phi_{\rm{o}_{1}}+(i-\rm{o}_{1})\epsilon_{\rm{o}},~{}\emph{i}\in[{\rm{o}_{1}},{\rm{o}_{2}}]&\\ \phi_{\rm{o}_{2}}+(i-\rm{o}_{2})\epsilon_{\rm{g}},~{}\emph{i}\in(\rm{o}_{2},\emph{D}]&.\end{array}\right.$ (4)
@@
Based on $\mathbf{A}$, the antenna-domain channel can be transformed into angular domain by $\mathbf{H}^{\rm{a}}_{k}=\mathbf{A}^{\rm H}\mathbf{H}_{k}$, where $\mathbf{H}^{\rm{a}}_{k}\in\mathbb{C}^{D\times K_{\rm{P}}}$ is the angular-domain channel between the $k$-th user and BS of the pilot position. Using $(\mathbf{A}^{\rm H})^{\dagger}\mathbf{A}^{\rm H}=\mathbf{I}_{N_{\rm t}}$, the received pilot signals can be rewritten as222 $\mathbf{A}$ is a row full rank matrix. $(\mathbf{A}^{\rm H})^{\dagger}=\mathbf{A}(\mathbf{A}^{\rm H}\mathbf{A})^{-1}$.:
$\mathbf{Y}^{\rm{P}}_{k}=\underbrace{\mathbf{F}_{\rm{R}}^{\rm H}(\mathbf{A}^{\rm H})^{\dagger}}_{\Theta}\mathbf{H}^{\rm{a}}_{k}\mathbf{S}^{\rm{P}}_{k}+\mathbf{F}_{\rm{R}}^{\rm H}\mathbf{N}^{\rm{P}}_{k}.$ (5)
Here, $\Theta=\mathbf{F}_{\rm{R}}^{\rm H}(\mathbf{A}^{\rm H})^{\dagger}$ is defined as the sensing matrix. Then, the pilot observations $\mathbf{Y}^{\rm{P}}_{k}$ are projected into the sensing matrix space by $\tilde{\mathbf{Y}}^{\rm{P}}_{k}=\Theta^{\dagger}\mathbf{Y}^{\rm{P}}_{k}$, which captures user-localized angular features. The complex value $\tilde{\mathbf{Y}}^{\rm{P}}_{k}\in\mathbb{C}^{D\times N_{\rm{p}}}$ is then transformed into real value input $\tilde{\mathbf{Y}}^{\rm{r}}_{k}\in\mathbb{R}^{2\times D\times N_{\rm{p}}}$ by treating the imaginary and real parts as two parallel dimensions. $\tilde{\mathbf{Y}}^{\rm{r}}_{k}$ serves as the input for a customized two-layer CNN, referred to as the pilot-based CSI feature extraction CNN (PCF-CNN).
@@ III-C OOB CSI Feature Extraction via Fine LiDAR Processing
In Eq. (6), $f=f_{\rm{up}}+f_{\rm{down}}$ is the vertical field of view (FoV) of LiDAR; $f_{\rm{up}}$ and $f_{\rm{down}}$ are the upward and downward FoV, respectively. The width ($w$) and height ($h$) of range image are determined by the horizontal angular resolution $a_{\rm{h}}$ and the number of beams $M_{\rm{b}}$ of the LiDAR, such that $w={360^{\circ}}/{a_{\rm{h}}}$ and $h=M_{\rm{b}}$.
Step 2: Multiple Signal Propagation (SP)-Feature Channels Superposition. Based on the LiDAR range image $\mathbf{R}_{\rm{D}}$, we propose to construct a multi-SP-feature superposed range image $\mathbf{R}_{\rm{D}}^{\rm{RF}}\in\mathbb{R}^{3\times w\times h}$. The first channel represents the original range image, denoted as $\mathbf{R}_{\rm{D}}^{\rm{RF}}[1,u_{i},v_{i}]=r_{i}$ where $r_{i}$ is the range at the pixel $\mathbf{r}_{i}$. The second channel is referred to as the equivalent small-scale fading map, which aims to provide information about potential multipath components. The pixel values in this channel, $\mathbf{R}_{\rm{D}}^{\rm{RF}}[2,u_{i},v_{i}]$, can take one of three possible values: receiver, vehicles near the receiver (potential scatterers), and other objects. The third channel is known as the equivalent large-scale fading map, designed to offer insights into signal amplitude variation characteristics. The pixel values in this channel, $\mathbf{R}_{\rm{D}}^{\rm{RF}}[3,u_{i},v_{i}]$, represent the path loss of the signal transmitted from transmitter to the point $\mathbf{r}_{i}$. The pixel values of the second and third channels are determined through specific criteria and considerations related to the environmental detection results acquired from LiDAR, and are expressed as:
$\mathbf{R}_{\rm{D}}^{\rm{RF}}[2,u_{i},v_{i}]=\left\{\begin{array}[]{cl}2,{}&\mathbf{p}_{i}\in\mathcal{V}_{\rm{r}}\\ 1,{}&\mathbf{p}_{i}\in\mathcal{V}_{\rm{s}}\\ 0,{}&\text{otherwise}\end{array}\right.$ (7a)
@@ III-D Feature Fusion via Attention
Although OOB wireless features closely related to CSI have been extracted from LiDAR and fine-grained user-localized CSI features have been extracted from pilot observations, a key challenge that remains to be addressed is how to intelligently regulate the utilization of both modalities under varying channel conditions. Drawing inspiration from the SENet channel attention mechanism [17], we have designed an adaptive feature weight control (AFWC) module to facilitate element-wise weight allocation between CSI features from the two modalities. Denoting the CSI features extracted from pilots and LiDAR as $\mathbf{q}_{\rm{P}}$ and $\mathbf{q}_{\rm{L}}$, respectively, they are concatenated and input into the AFWC module, which learns the importance of different features and outputs the weight vector $\mathbf{w}$:
@@
Here, $\varphi_{\rm w}^{l}(\cdot)=f_{\rm{ReLU}}(\mathbf{W}_{\rm w}^{l}\mathbf{X}+\mathbf{b}_{\rm w}^{l}),l=1,2$ is the non-linear functions of the first and second fully connected (FC) layers with ReLU as the activation function; $\varphi_{\rm{w}}^{3}(\cdot)=f_{\rm{Sigmoid}}(\mathbf{W}_{\rm w}^{3}\mathbf{X}+\mathbf{b}_{\rm w}^{3})$ is the non-linear function of the third layer with Sigmoid as the activation function; $\oplus$ represents the concatenation operation. $\mathbf{X}$ denotes a certain input, while $\mathbf{W}_{\rm w}^{l}$ and $\mathbf{b}_{\rm w}^{l}$ denote the weight and bias of the $l$-th layer, respectively. Next, a weighted feature $\mathbf{q}^{\rm{w}}$ is formed by performing an element-wise product between the weight vector $\mathbf{w}$ and the raw multi-modal CSI feature $\mathbf{q}=\mathbf{q_{\rm P}}\oplus\mathbf{q_{\rm L}}$, denoted as $\mathbf{q}^{\rm{w}}=\mathbf{q}\otimes\mathbf{w}$.
@@ III-E CSI Reconstruction
$\hat{\mathbf{H}}_{k}^{\rm{P}}=\text{MLP}_{\rm P}(\mathbf{q}^{\rm{w}})=\varphi_{\rm P}^{L_{\rm P}}(\varphi_{\rm P}^{L_{\rm P}-1}(\cdots\varphi_{\rm P}^{1}(\mathbf{q}^{\rm{w}})\cdots)).$ (9)